A box of mass m that was sliding along the floor, hits an incline plane and slides up the inclined plane with angle theta at a velocity v0. Find an expression for the maximum height, h, above the floor that the box reaches. This expression should not depend on m or theta, but only on v0 and g.
Answer #1

assuming no frictional force is acting, by conservation of energy

KE+PE = constant

at ground = only KE = 1/2mv0^2

at highest point on inclined = only PE = mgh

mgh = 1/2mv0^2

h = v0^2/2g
Answer #2


a =gsintheta

Vo *Vo = 2*gsintheta * S

S = (Vo^2)/(2gsintheta) ( S distance along plane )

h= vertical heigh from ground

Ssintheta =h

h = (Vo^2)/(2g)